40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 16:56:38 2013, Jenne, Update, LSC, LSC model sensing matrix upgrades in progress 
    Reply  Wed Oct 9 17:27:35 2013, Jenne, Update, LSC, LSC model sensing matrix upgrades  
       Reply  Wed Oct 9 22:05:55 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
          Reply  Fri Oct 11 17:37:08 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
             Reply  Sun Oct 13 13:29:09 2013, rana, Update, LSC, New lockin / sensing matrix screens 
                Reply  Mon Oct 14 17:40:15 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
                   Reply  Thu Oct 17 13:40:55 2013, Jenne, Update, LSC, New lockin / sensing matrix frequencies All_REFL_PRMI_17Oct2013.pdf
             Reply  Mon Oct 14 21:12:35 2013, Jenne, Update, LSC, New lockin / sensing matrix screens 
          Reply  Wed Oct 16 03:04:37 2013, Jenne, Update, LSC, New lockin / sensing matrix model parts 
             Reply  Wed Oct 23 19:28:22 2013, Jenne, Update, LSC, New lockin / sensing matrix model parts 
Message ID: 9234     Entry time: Fri Oct 11 17:37:08 2013     In reply to: 9227     Reply to this: 9236   9239
Author: Jenne 
Type: Update 
Category: LSC 
Subject: New lockin / sensing matrix screens 

Quote:

Still to do:

* Decide what 5 frequencies to use for excitation.

* Put the bandpass and lowpass filters into the lockin modules.

* Put a little more stuff into the front end so that we get total mag and phase of the sensing matrix element, not just uncalibrated lockin outputs.

 I worked on some of these "to-do" things today for the new sensing matrix setup.  I chose several frequencies around 90 Hz for my measurements.  This was an area (while PRMI was locked with REFL 165 I&Q, and all 4 REFL PDs had their whitening on) there was a pretty wide clean space in the noise spectrum.

I put bandpass filters into the _SIG banks of the lockin modules, and lowpass filters into the _I banks.  However, when I loaded the new filter coefficients, it looks like not all of them are showing up in the screens.  I'm a little confused as to why.  They are still there if I close and re-open Foton, so I think I really put them in.

Also, I don't think that I'm successfully getting any signal from the LSC model into the new lockin modules on the CAL model.  I'm not sure if this is to do with the fact that I added 32 more IPC connections the other day.  I've emailed Jamie to ask whether or not we may have hit some limit.

I also tried testing out a small bit of c-code for the calibration of the lockin outputs.  It seems as though I cannot do an arctangent in the front end.  When I compile the c1tst model, and start it up, if I have an "atan2" in the code, it tells me "No Sync".  However, if I remove that line of c-code, the model compiles fine (which it did in the case with the arctan), and the model runs just fine (which it doesn't with the arctan).  My backup plan is to include a Taylor expansion for the arctangent in some c-code.

ELOG V3.1.3-